# Teleoscope: Exploring Themes in Large Document Sets By Example

### v1

@@ 5. Discussion, Limitations and Future Work
Through our studies and design work, we settled on a process-focused graph-based workflow metaphor rather than a dashboard metaphor (DG2). This was due to our focus on the process of arranging and connecting documents as being the most important interaction focus (DG1, DG3); it also increased visibility of system state and allowed for direct inspection of computational elements. Researchers found that they were able to find confidence that they had searched “enough” of the document space through using the Teleoscope ranking operation, and found potential in the Projection clustering operation in supporting rigorous document space exploration (E2.RQ2, E3.RQ2). After we iterated on our design to allow for collaboration and make it obvious to our participants how to use collaboration features, they were able to satisfy their need for theoretical saturation through collaborative iterative searches using Teleoscope (E2.RQ1, DG2).
@@
A limitation of a workflow metaphor is that it does not visually model the unknown document space very well. In part, this helps manage the cognitive overload since users are only shown what they have explicitly requested and built. However, having some dashboard elements such as summary statistics of corpus exploration relative to selected groups would be an obvious next design step. In our estimation, this would require a more sophisticated graph processing approach.
@@
Due to our users’ use of an external keyword Google doc, an open question is how and whether to incorporate keywords back into Teleoscope as a primary interaction element (E3.RQ1). Perhaps with more experience with Teleoscope, set operations would be used to manage keywords more directly. However, it may not be necessary to treat keywords as a primary interaction element within Teleoscope. We already imagine Teleoscope as part of a qualitative research tool ecosystem: current research shows that Teleoscope can work within existing analysis workflows between tools such as NVivo or PowerBI. We found it interesting and useful that researchers were able to creatively include Teleoscope in their research process using external tools that they were already familiar with, and to iterate between them and Teleoscope. For this reason, we support exporting to common document formats such as XLSX, DocX and JSON. However, having interface standards that make Teleoscope output interoperate with other qualitative software is important.
Since this is a research product, Teleoscope does not yet incorporate standard enterprise product features such as arbitrary data import. We imagine that an enterprise version of Teleoscope would need to support swapping/encoding with different exogenous models directly in the user interface, particularly as large language models become more sophisticated.
@@
In our estimation, a large part of Teleoscope’s success is demonstrated by its on-going public release and use by the research team that we recruited for field deployment in Study 3. We are continuing to recruit more research teams and hope to see Teleoscope develop into a more mature product as they use it.
@@
Teleoscope’s impact on methodology. To what extent is it desirable for an ML model to “think like you” when you are a qualitative researcher? One of the most interesting and relevant questions for Teleoscope is the way in which it can impact qualitative research methodologies. This was brought up by multiple participants, and is an important question for both adoption of Teleoscope and for shedding light on other burgeoning ML-supported methodologies that deal with data interpretation.
@@
Current qualitative methodologies place a high level of importance on disciplined and ethical data collection, analysis, and reporting. Since interpretation necessarily comes from the perspective of the researcher, researchers attempt to maintain distance from analysis until data collection is concluded. Teleoscope could be seen as presupposing a thematic structure before detailed analysis has begun, which would violate the principle of letting the themes emerge from the data.
@@
However, our participants articulated this mostly as fear of cognitive bias about the data, not that it fundamentally interfered with having an opinionated approach to data collection. It is impossible to do data collection from a truly unbiased perspective, since researchers still have editorial discretion in choosing their research interests, topics, and data sources. Our participants were satisfied with Teleoscope bringing up unexpected results along with expected results as a way to counteract bias. So, the answer to the question at the start of this section is that you might want the ML model to “think like you” when you are searching for expected results, and then to not “think like you” when you are challenging your expectation. Teleoscope can do both.
@@
We believe that Teleoscope provides an opportunity to directly inspect and compare biases by inspecting and comparing the workflows that trace how a document collection came to be. Returning to the Nowell quote from Section 2, we believe that Teleoscope can demonstrate that data collection has been done in a precise, consistent, and exhaustive manner through tracing the data collection process with enough detail to enable the research team to determine whether the process is credible (DG1, DG2).
@@ 6. Conclusion
In this paper, we presented Teleoscope, a web-based system that supports interactive exploration of large corpora (100K-1M) of short documents. We developed it in response to the need of qualitative researchers to explore large corpora in meaning-based ways using natural interaction techniques. Targeting qualitative “themes” rather than “topics,” we attempted to create a system that makes an ML model “think like you” without significant retraining. Teleoscope provides ML-based workflows that have semantic and computational meaning. These workflows help researchers to retrace, share, and recompute their sense-making process. We reported on the design, engineering, evaluation, and deployment of our system. Our public deployment of Teleoscope is ongoing. We plan to continue improving on Teleoscope and to maintain it for use by the broad community of qualitative researchers.

### v2

@@ 6. Design Choices and Limitations
Through our studies and design work, we settled on a process-focused graph-based workflow metaphor rather than a dashboard metaphor (DG2). This was due to our focus on the process of arranging and connecting documents as being the most important interaction focus (DG1, DG3); it also increased visibility of system state and allowed for direct inspection of computational elements. Researchers found that they were able to find confidence that they had searched “enough” of the document space through using the Ranking operation, and found potential in the Projection clustering operation in supporting rigorous document space exploration (CS2.RQ2). After we iterated on our design to allow for collaboration and make it obvious to our participants how to use collaboration features, they were able to satisfy their need for theoretical saturation/information power through collaborative iterative searches using Teleoscope (CS2.RQ1, DG2).
@@
A limitation of a workflow metaphor is that it does not visually model the unknown document space very well. In part, this helps manage the cognitive overload since users are only shown what they have explicitly searched for themselves. However, some dashboard elements such as summary statistics of corpus exploration relative to selected groups would be an obvious next design step.
@@
Due to our users’ use of an external keyword Google doc, an open question is how and whether to incorporate keywords back into Teleoscope as a primary interaction element (CS2.RQ1). Perhaps with more experience with Teleoscope, set operations would be used to manage keywords more directly. However, we do not imagine Teleoscope as an end-to-end analysis tool; instead, we imagine Teleoscope as part of a qualitative research tool ecosystem. Teleoscope can work within existing analysis workflows between tools such as NVivo or PowerBI using our import/export feature to common document formats such as XLSX, DocX and JSON. Researchers were able to creatively include Teleoscope using external tools and to iterate between them and Teleoscope.
@@
In our opinion, a large part of Teleoscope’s success is demonstrated by its on-going public release and use by the research team that we recruited for field deployment in Case Study 3. We are continuing to recruit more research teams and hope to see Teleoscope develop into a more mature product as they use it.
@@
Teleoscope’s impact on methodology. One of the most interesting and relevant questions for Teleoscope is the way in which it can impact qualitative research methodologies. This was brought up by multiple participants, and is an important question for both adoption of Teleoscope and for shedding light on other burgeoning ML-supported methodologies that deal with data interpretation.
@@
Current qualitative methodologies place a high level of importance on disciplined and ethical data curation, analysis, and reporting. Since interpretation necessarily comes from the perspective of the researcher, researchers attempt to maintain distance from analysis until data curation is concluded. Teleoscope could be seen as presupposing a thematic structure before detailed analysis has begun, which would violate the principle of letting the themes emerge from the data.
@@
However, our participants articulated this mostly as fear of cognitive bias about the data, not that it fundamentally interfered with having an opinionated approach to data curation. It is impossible to do data curation from a truly unbiased perspective, since researchers still have editorial discretion in choosing their research interests, topics, and data sources. Our participants were satisfied with Teleoscope bringing up unexpected results along with expected results as a way to counteract bias.
@@
We believe that Teleoscope provides an opportunity to directly inspect and compare biases by inspecting and comparing the workflows that trace how a document collection came to be. Returning to the Nowell quote from Section 2, we believe that Teleoscope can demonstrate that data collection has been done in a precise, consistent, and exhaustive manner through tracing the data curation process with enough detail to enable the research team to determine whether the process is credible (DG1, DG2).
@@ 7. Conclusion
In this paper, we presented Teleoscope, a web-based system that supports interactive exploration of large corpora (100K-1M) of short documents. We developed it in response to the need of qualitative researchers to explore large corpora in meaning-based ways using natural interaction techniques. Teleoscope provides ML-based workflows that have semantic and computational meaning. These workflows help researchers to retrace, share, and recompute their sense-making process. We reported on the design, engineering, evaluation, and deployment of our system. Our public deployment of Teleoscope is ongoing (Teleoscope.ca, 2024a). We plan to continue improving on Teleoscope and to maintain it for use by the broad community of qualitative researchers.